OpenStudy (anonymous):

If f(x) = −5x + 1, find f(−3). 15 −14 −7 16

OpenStudy (anonymous):

The answer is 16. :)

OpenStudy (imtiaz7):

f(x)= -5x+1 put the value of x as -3: f(-3)= -5(-3)+1 f(-3)= 15+1 f(-3)= 16